# MaskGCT: Zero-Shot Text-to-Speech with Masked Generative Codec Transformer

### v1

@@ Appendix C Details of Duration Predictor
MaskGCT requires specifying the target speech duration during inference, so we train a flow matching [Lipman et al., 2022, Liu et al., 2022b] based duration predictor to obtain the total duration of the target audio by summing the phone-level duration. Note that we do not need to actually use the phone-level durations but only use them to make a reasonable estimate of the total duration, leaving other total duration predictor methods for future works to explore. The duration predictor has a similar Transformer architecture to MaskGCT, with 12 layers, 12 attention heads, and a hidden size of 768. We also adapt in-context learning and classifier-free guidance for the duration predictor. During training, we randomly select a prefix segment of the phoneme sequence and its corresponding duration as a prompt, which is not added with noise. At the same time, we use a probability of 0.15 to drop the prompt. We model the duration in the log domain using flow matching. We denote $x_{1}$ as a random variable of $\log(\text{duration}+1)$, $x_{0}$ as a randomly sampled Gaussian noise, then $v_{\theta}(x_{t},t)=x_{t}=(1-t)x_{0}+tx_{1}$, where the timestep $t\in[0,1]$. The loss function of the duration predictor is $\mathbb{E}_{t,x_{1}}(v_{\theta}(x_{t},t)-(x_{1}-x_{0}))^{2}$. In the inference stage, we use a midpoint ODE solver to generate the target from randomly sampled Gaussian noise with a total of 4 steps. We pretrain a duration aligner (between phoneme and W2v-BERT 2.0 semantic feature) based on monotonic alignment search (MAS) [Kim et al., 2020] to get the ground truth duration for each phoneme.
Appendix D Text Tokenizer
@@
We consider two text tokenization methods: Grapheme-to-Phoneme (G2P) and Byte Pair Encoding (BPE). For G2P, we employ phonemize  [^3] https://github.com/bootphon/phonemizer for English and a combination of jieba  [^4] https://github.com/fxsjy/jieba and pypinyin  [^5] https://github.com/mozillazg/python-pinyin for Chinese. For BPE, we utilize the BPE method and vocabulary from Whisper  [^6] https://github.com/huggingface/transformers/blob/main/src/transformers/models/whisper/tokenization_whisper.py, with a vocabulary size exceeding 30,000.
Appendix E Classifier-Free Guidance
@@
We adopt the classifier-free guidance [Ho and Salimans, 2022] technique for both the T2S model and the S2A model. We also introduce classifier-free guidance with rescaling, following Lin et al. [2024]. In the training stage, we randomly drop the prompt with a probability of 0.15 to model the probability distribution $p_{\theta}(\mathbf{X})$ without the prompt. During inference, we compute the output embedding $g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g_{\theta}(\mathbf{X}|\mathbf{X}^{p})+w_{\text{cfg}}\cdot(g_{\theta}(\mathbf{X}|\mathbf{X}^{p})-g_{\theta}(\mathbf{X}))$ of the last layer of the model, where $w_{\text{cfg}}$ is the classifier-free guidance scale, then we compute the rescale embedding $g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})\times\text{std}(g_{\theta}(\mathbf{X}|\mathbf{X}^{p}))/\text{std}(g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p}))$, the final output embedding is computed as $w_{\text{rescale}}\cdot g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})+(1-w_{\text{rescale}})\cdot g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})$. In our paper, $w_{\text{cfg}}$ and $w_{\text{rescale}}$ are set as 2.5 and 0.75 by default.
Appendix F Evaluation Baselines
@@
VALL-E [Wang et al., 2023]. A large-scale TTS system uses an autoregressive and an additional non-autoregressive model to predict discrete tokens from a neural speech codec [Défossez et al., 2022]. We reproduce VALL-E with Amphion toolkit [Zhang et al., 2023] and Librilight [Kahn et al., 2020] dataset.
@@
NaturalSpeech 3 [Ju et al., 2024]. A non-autoregressive model large-scale TTS systems with factorized speech codec for speech decoupling representation and factorized diffusion Models for speech generation. It achieves human-level naturalness on the LibriSpeech test set. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
@@
VoiceBox [Le et al., 2024]. A non-autoregressive model large-scale multi-task speech generation model based on flow matching Lipman et al. [2022]. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
@@
XTTS-v2 [Casanova et al., 2024]. An open-source multilingual TTS model that supports 16 languages. It is also based on an autoregressive model. We use the official code and pre-trained checkpoint  [^7] https://huggingface.co/coqui/XTTS-v2.
@@
VoiceCraft [Peng et al., 2024]. A token-infilling neural codec language model for text editing and text-to-speech. It predicts multi-layer tokens in a delay pattern. We use the official code and pre-trained checkpoint  [^8] https://huggingface.co/pyp1/VoiceCraft/blob/main/830M_TTSEnhanced.pth.
@@
CosyVoice [Du et al., 2024a]. A two-stage large-scale TTS system. The first stage is an autoregressive model and the second stage is a diffusion model. It is trained on 170,000 hours of multilingual speech data. We use the official code and pre-trained checkpoint  [^9] https://huggingface.co/model-scope/CosyVoice-300M.
@@ Appendix G Post-Training for Emotion Control
MaskGCT can unlock more extensive capabilities with post-training. We take emotion control as an example. After being pretrained on a large-scale dataset, we fine-tune the T2S model by adding an additional emotion label as a prefix to the original input sequence. We use a small emotion dataset, ESD [Zhou et al., 2021], which consists of 350 parallel utterances with an average duration of 2.9 seconds spoken by 10 native English and 10 native Mandarin speakers, to fine-tune our model. The experimental results show that MaskGCT can unlock emotion control capabilities for zero-shot in-context learning scenarios. For the construction of the train and test datasets, we selected one male and one female speaker each from native English and native Mandarin backgrounds, resulting in a total of four speakers for the test dataset. The remaining 16 speakers were allocated to the train dataset. For the 350 parallel Chinese utterances, we randomly chose 22 utterances for the test set, with the remaining utterances designated for training. Similarly, for the 350 parallel English utterances, we randomly selected 21 utterances for the test set, with the rest used for training. To assess the consistency between the generated audio and the target emotion label, we trained an emotion classification model using the constructed train dataset. This model achieved a classification accuracy of 72% on the test dataset. We show some examples in our demo page.
Appendix H Speech Editing
Based on the mask-and-predict mechanism, our text-to-semantic model supports zero-shot speech content editing with the assistance of a text-speech aligner. By using the aligner, we can identify the editing boundary of the original semantic token sequence, mask the portion that needs to be edited, and then predict the masked semantic tokens using the edited text and the unmasked semantic tokens. However, we have observed that our system is not very robust in editing tasks. A possible conjecture is that we need to adopt a training paradigm better suited for editing tasks, such as fill-in-mask [Le et al., 2024, Du et al., 2024b]. We show some examples in our demo page.
@@ Appendix I Voice Conversion
MaskGCT supports zero-shot voice conversion by fine-tuning the S2A with a modified training strategy. The zero-shot voice conversion task aims to alter the source speech to sound like that of a target speaker using a reference speech from the target speaker, without changing the semantic content. We can directly use the semantic tokens $\mathbf{S}_{\text{src}}$ extracted from the source speech and the prompt acoustic tokens $\mathbf{A}_{\text{ref}}$ extracted from the reference speech to predict the target acoustic tokens $\mathbf{A}_{\text{tgt}}$. Since $\mathbf{S}_{\text{src}}$ may retain some timbre information, we perform timbral perturbation on the semantic features input to the semantic codec encoder. Specifically, we apply timbral perturbation to the input mel-spectrogram features of the W2v-BERT 2.0 model, following the method outlined in FreeVC [Li et al., 2023b]. We fine-tune our S2A model using this training strategy. We show some examples in our demo page.
Appendix J Duration Controllable Speech Translation
The goal of the speech translation task is to translate speech from one language to another while preserving the original semantic, timbre, and prosody. In some scenarios, we also need to ensure that the total duration remains relatively unchanged, such as in video translation. Our model can achieve this seamlessly, with the ability to control the total duration and, through in-context learning, use the pre-translation speech as a prompt to maintain the timbre and prosody. We show some examples of speech translation between Chinese and English in our demo page.
Appendix K Hard Cases Evaluation
@@
Appendix L Discussion about Concurrent Works
SimpleSpeech [Yang et al., 2024], DiTTo-TTS [Lee et al., 2024], and E2 TTS [Eskimez et al., 2024] are also NAR-based models that do not necessitate precise alignment information between text and speech, nor do they forecast phoneme-level duration. These are concurrent works with MaskGCT. The three models all employ diffusion modeling on speech representations within continuous spaces. SimpleSpeech models the latent representation of a wav codec based on finite scalar quantization (FSQ) [Mentzer et al., 2023], DiTTo-TTS utilizes the latent representation of a wav codec based on residual vector quantization (RVQ), and E2 TTS directly models the mel-spectrogram with flow matching.
Appendix M Boarder Impact

### v2

@@ A.5 Details of Duration Predictor
MaskGCT requires specifying the target speech duration during inference, so we train a flow matching Lipman et al. [2022], Liu et al. [2022b] based duration predictor to obtain the total duration of the target audio by summing the phone-level duration. Note that we do not need to actually use the phone-level durations but only use them to make a reasonable estimate of the total duration, leaving other total duration predictor methods for future works to explore. The duration predictor has a similar Transformer architecture to MaskGCT, with 12 layers, 12 attention heads, and a hidden size of 768. We also adapt in-context learning and classifier-free guidance for the duration predictor. During training, we randomly select a prefix segment of the phoneme sequence and its corresponding duration as a prompt, which is not added with noise. At the same time, we use a probability of 0.15 to drop the prompt. We model the duration in the log domain using flow matching. We denote $x_{1}$ as a random variable of $\log(\text{duration}+1)$, $x_{0}$ as a randomly sampled Gaussian noise, then $v_{\theta}(x_{t},t)=x_{t}=(1-t)x_{0}+tx_{1}$, where the timestep $t\in[0,1]$. The loss function of the duration predictor is $\mathbb{E}_{t,x_{1}}(v_{\theta}(x_{t},t)-(x_{1}-x_{0}))^{2}$. In the inference stage, we use a midpoint ODE solver to generate the target from randomly sampled Gaussian noise with a total of 4 steps. We pretrain a duration aligner (between phoneme and W2v-BERT 2.0 semantic feature) based on monotonic alignment search (MAS) Kim et al. [2020] to get the ground truth duration for each phoneme.
A.6 Text Tokenizer
Table 10: G2P vs. BPE.
SIM-O $\uparrow$ WER $\downarrow$
SeedTTS test-en
G2P 0.728 2.466
BPE 0.711 4.036
SeedTTS test-zh
G2P 0.777 2.183
BPE 0.769 1.921
@@ A.6 Text Tokenizer
We consider two text tokenization methods: Grapheme-to-Phoneme (G2P) and Byte Pair Encoding (BPE). For G2P, we employ phonemize [^6] https://github.com/bootphon/phonemizer for English and a combination of jieba [^7] https://github.com/fxsjy/jieba and pypinyin [^8] https://github.com/mozillazg/python-pinyin for Chinese. For BPE, we utilize the BPE method and vocabulary from Whisper [^9] https://github.com/huggingface/transformers/blob/main/src/transformers/models/whisper/tokenization_whisper.py, with a vocabulary size exceeding 30,000. Table 10 shows the comparison results of MaskGCT using the two different text tokenization methods. The results indicate that G2P outperforms BPE in English with a higher SIM-O of 0.728 compared to 0.711 and a lower WER of 2.466 versus 4.036. Conversely, in Chinese, G2P maintains a slightly higher SIM-O (0.777 vs. 0.769) but BPE achieves a lower WER (1.921 vs. 2.338). These findings suggest that while G2P is superior in preserving text similarity and reducing errors in English, BPE is more effective in minimizing WER in Chinese. We hypothesize that the reason might be that the Chinese G2P system we used still has deficiencies in handling polyphonic characters. In contrast, BPE can learn different pronunciations for the same character based on context.
Appendix B Discussion about Semantic and Acoustic Definitions
In this paper, we refer to the speech representation extracted from the speech self-supervised learning (SSL) model as the semantic feature. The discrete tokens obtained through the discretization of these semantic features (using k-means or vector quantization are termed semantic tokens. Similarly, we define the representations from melspectrogram, neural speech codecs, or speech VAE as acoustic features, and their discrete counterparts are called acoustic tokens. This terminology was first introduced in Borsos et al. [2023b] and has since been adopted by many subsequent works Borsos et al. [2023a], Guo et al. [2024], Ju et al. [2024], Huang et al. [2023], Zhang et al. [2023a]. It is important to note that this is not a strictly rigorous definition. Generally, we consider semantic features or tokens to contain more prominent linguistic information and exhibit stronger correlations with phonemes or text. One measure of this is the phonetic discriminability in terms of the ABX error rate. In this paper, the W2v-BERT 2.0 features we use have a phonetic discriminability within less than 5 on the LibriSpeech dev-clean dataset, whereas acoustic features, for example, Encodec latent features, score above 20 on this metric. However, it is worth noting that semantic features or tokens not only contain semantic information but also include prosodic and timbre aspects. In fact, we suggest that for certain two-stage zero-shot TTS systems, excessive loss of information in semantic tokens can degrade the performance of the second stage, where semantic-to-acoustic conversion occurs. Therefore, finding a speech representation that is more suitable for speech generation remains a challenging problem.
Appendix C Classifier-Free Guidance
@@
We adopt the classifier-free guidance Ho and Salimans [2022] technique for both the T2S model and the S2A model. We also introduce classifier-free guidance with rescaling, following Lin et al. [2024]. In the training stage, we randomly drop the prompt with a probability of 0.15 to model the probability distribution $p_{\theta}(\mathbf{X})$ without the prompt. During inference, we compute the output embedding $g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g_{\theta}(\mathbf{X}|\mathbf{X}^{p})+w_{\text{cfg}}\cdot(g_{\theta}(\mathbf{X}|\mathbf{X}^{p})-g_{\theta}(\mathbf{X}))$ of the last layer of the model, where $w_{\text{cfg}}$ is the classifier-free guidance scale, then we compute the rescale embedding $g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})\times\text{std}(g_{\theta}(\mathbf{X}|\mathbf{X}^{p}))/\text{std}(g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p}))$, the final output embedding is computed as $w_{\text{rescale}}\cdot g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})+(1-w_{\text{rescale}})\cdot g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})$. In our paper, $w_{\text{cfg}}$ and $w_{\text{rescale}}$ are set as 2.5 and 0.75 by default.
Appendix D Evaluation Baselines
@@
VALL-E Wang et al. [2023]. A large-scale TTS system uses an autoregressive and an additional non-autoregressive model to predict discrete tokens from a neural speech codec Défossez et al. [2022]. We reproduce VALL-E with Amphion toolkit Zhang et al. [2023b] and Librilight Kahn et al. [2020] dataset.
@@
NaturalSpeech 3 Ju et al. [2024]. A non-autoregressive model large-scale TTS systems with factorized speech codec for speech decoupling representation and factorized diffusion Models for speech generation. It achieves human-level naturalness on the LibriSpeech test set. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
@@
VoiceBox Le et al. [2024]. A non-autoregressive model large-scale multi-task speech generation model based on flow matching Lipman et al. [2022]. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
@@
XTTS-v2 Casanova et al. [2024]. An open-source multilingual TTS model that supports 16 languages. It is also based on an autoregressive model. We use the official code and pre-trained checkpoint [^10] https://huggingface.co/coqui/XTTS-v2.
@@
VoiceCraft Peng et al. [2024]. A token-infilling neural codec language model for text editing and text-to-speech. It predicts multi-layer tokens in a delay pattern. We use the official code and pre-trained checkpoint [^11] https://huggingface.co/pyp1/VoiceCraft/blob/main/830M_TTSEnhanced.pth.
@@
CosyVoice Du et al. [2024a]. A two-stage large-scale TTS system. The first stage is an autoregressive model and the second stage is a diffusion model. It is trained on 170,000 hours of multilingual speech data. We use the official code and pre-trained checkpoint [^12] https://huggingface.co/model-scope/CosyVoice-300M.
Appendix E Multilingual Zero-Shot TTS
We validate the effectiveness of MaskGCT across four additional languages beyond Chinese and English, specifically Japanese, Korean, German, and French. On the foundation of our existing training data, we expand by 2,500 hours of Japanese, 7,400 hours of Korean, 6,900 hours of German, and 8,200 hours of French. We collect these data using the data collection pipeline proposed by He et al. [2024]. For evaluation, we use the test sets provided in He et al. [2024]. We still employ SIM-O and WER as evaluation metrics, with Whisper-medium131313ttps://huggingface.co/openai/whisper-medium serving as the ASR model for WER assessment. We utilize XTTS-v2 and the two models proposed in He et al. [2024]: Emilia-AR and Emilia-NAR as comparative baselines. Table 11 shows the results. MaskGCT demonstrates significant improvements over the baselines, with the exception of WER in Japanese. It is noteworthy that we only retrained our text-to-semantic model using the expanded data, without retraining the tokenizers and semantic-to-acoustic models. We believe that further enhancements in our model’s performance can be achieved if all components are retrained on the expanded data.
Table 11: Evaluation results for MaskGCT and baseline methods on the test sets for Japanese, Korean, German, and French.
System Ja Ko Fr De
WER SIM-O WER SIM-O WER SIM-O WER SIM-O
Emilia-AR 3.6 0.625 10.9 0.681 8.2 0.589 6.8 0.680
Emilia-NAR 10.8 0.562 15.2 0.608 17.5 0.550 13.3 0.633
XTTS-v2 2.981 0.579 12.45 0.617 6.898 0.531 9.168 0.569
MaskGCT 3.903 0.678 9.417 0.732 5.598 0.667 5.126 0.745
Appendix F Duration-Controllable Speech Translation
The goal of the speech translation task is to translate speech from one language to another while preserving the original semantic, timbre, and prosody. In some scenarios, we also need to ensure that the total duration remains relatively unchanged, such as in cross-lingual dubbing. Our model can achieve this seamlessly, with the ability to control the total duration and, through in-context learning, use the pre-translation speech as a prompt to maintain the timbre and prosody. To quantify the capabilities of our model, we randomly select 200 samples from SeedTTS test-zh and 200 samples from SeedTTS test-en. Additionally, we sample 200 examples for each language of Japanese, Korean, German, and French from each of the test sets provided in He et al. [2024]. Subsequently, we utilize GPT4o-mini Achiam et al. [2023] to translate each sample into one of the other five languages, using the translated text as the target text. We use the duration of prompt speech as the duration of target speech. This process yields 30 sets of test data. Table 12 shows the results of the 30 sets of experiments. We observe that MaskGCT maintains a good level of speaker similarity across translations between the six languages. Both “X to En” and “En to X” generally perform well, characterized by relatively low WER values and moderate SIM-O scores. “X to Ja” also achieve low WER values. However, for languages other than English, “X to Zh”, “X to De”, and “X to Fr” exhibit higher WER values. We hypothesize that the primary reasons for this include the difficulty in maintaining accurate pronunciation while preserving the same duration before and after translation, as well as the limited training data for Fr and De. Achieving more robust cross-lingual translation remains a focus for future work. We also show some examples of speech translation in our demo page.
Table 12: Evaluation results in cross-lingual speech translation with consistent total duration.
Zh En Ja Ko De Fr
WER SIM-O WER SIM-O WER SIM-O WER SIM-O WER SIM-O WER SIM-O
Zh - - 7.466 0.678 7.864 0.720 9.751 0.736 25.54 0.724 16.21 0.687
En 7.411 0.535 - - 5.870 0.544 12.18 0.543 12.43 0.579 17.48 0.590
Ja 13.93 0.647 7.387 0.642 - - 10.98 0.703 12.85 0.649 14.61 0.645
Ko 31.30 0.734 14.61 0.697 12.79 0.749 - - 26.58 0.722 33.96 0.712
De 19.54 0.714 5.148 0.740 6.072 0.678 12.02 0.667 - - 14.53 0.672
Fr 32.84 0.672 12.17 0.682 6.076 0.640 12.07 0.582 21.65 0.682 - -
@@ Appendix G Post-Training for Emotion Control
MaskGCT can unlock more extensive capabilities with post-training. We take emotion control as an example. After being pretrained on a large-scale dataset, we fine-tune the T2S model by adding an additional emotion label as a prefix to the original input sequence. We use an emotion dataset, ESD Zhou et al. [2021], which consists of 350 parallel utterances with an average duration of 2.9 seconds spoken by 10 native English and 10 native Mandarin speakers, to fine-tune our model. The experimental results show that MaskGCT can unlock emotion control capabilities for zero-shot in-context learning scenarios. For the construction of the train and test datasets, we selected one male and one female speaker each from native English and native Mandarin backgrounds, resulting in a total of four speakers for the test dataset. The remaining 16 speakers were allocated to the training dataset. For the 350 parallel Chinese utterances, we randomly chose 22 utterances for the test set, with the remaining utterances designated for training. Similarly, for the 350 parallel English utterances, we randomly selected 21 utterances for the test set, with the rest used for training. To assess the consistency between the generated audio and the target emotion label, we trained an emotion classification model using the constructed train dataset. This model achieved a classification accuracy of 72% on the test dataset. We show some examples in our demo page.
Appendix H Speech Content Editing
Based on the mask-and-predict mechanism, our text-to-semantic model supports zero-shot speech content editing with the assistance of a text-speech aligner. By using the aligner, we can identify the editing boundary of the original semantic token sequence, mask the portion that needs to be edited, and then predict the masked semantic tokens using the edited text and the unmasked semantic tokens. However, we have observed that our system is not very robust in editing tasks. A possible conjecture is that we need to adopt a training paradigm better suited for editing tasks, such as fill-in-mask Le et al. [2024], Du et al. [2024b]. We show some examples in our demo page.
@@ Appendix I Voice Conversion
MaskGCT supports zero-shot voice conversion by fine-tuning the S2A with a modified training strategy. The zero-shot voice conversion task aims to alter the source speech to sound like that of a target speaker using a reference speech from the target speaker, without changing the semantic content. We can directly use the semantic tokens $\mathbf{S}_{\text{src}}$ extracted from the source speech and the prompt acoustic tokens $\mathbf{A}_{\text{ref}}$ extracted from the reference speech to predict the target acoustic tokens $\mathbf{A}_{\text{tgt}}$. Since $\mathbf{S}_{\text{src}}$ may retain some timbre information, we perform timbral perturbation on the semantic features input to the semantic codec encoder. Specifically, we apply timbral perturbation to the input mel-spectrogram features of the W2v-BERT 2.0 model, following the method outlined in FreeVC Li et al. [2023b]. We fine-tune our S2A model using this training strategy. We show some examples in our demo page.
Appendix J Hard Cases Evaluation
@@
Appendix K Discussion about Concurrent Works
SimpleSpeech Yang et al. [2024], DiTTo-TTS Lee et al. [2024], and E2 TTS Eskimez et al. [2024] are also NAR-based models that do not necessitate precise alignment information between text and speech, nor do they forecast phoneme-level duration. These are concurrent works with MaskGCT. The three models all employ diffusion modeling on speech representations within continuous spaces. SimpleSpeech models the latent representation of a wav codec based on finite scalar quantization (FSQ) Mentzer et al. [2023], DiTTo-TTS utilizes the latent representation of a wav codec based on residual vector quantization (RVQ), and E2 TTS directly models the mel-spectrogram with flow matching.
Appendix L Boarder Impact